# RLVR-World: Training World Models with Reinforcement Learning

### v1

@@ A.3 Model Predictive Control for Web Agents
Following the method introduced in WMA [8], we use a policy model to sample 20 candidate actions and select the three most frequently sampled for further evaluation. For each selected action, our trained world model performs next-state prediction. A summarization model is then used to (1) identify the top 10 most salient state changes in the predicted next state, and (2) generate a natural language summary describing those changes. This transition summary, along with the task objective, is provided to a value model, which assigns a score between 1 and 5. This scoring query is repeated 20 times, and the final score for each action is computed as the average of these 20 responses. The action with the highest average score is selected for execution. Except for the world model, all models, including policy, summarization, and value, are implemented by prompting DeepSeek-V3 with top-$p$ sampling ($p=0.95$). All prompts used in model predictive control are detailed in Appendix D.
@@ A.4 Robot Manipulation Trajectory Prediction
Hyperparameters of architectures and training process for robot manipulation trajectory prediction are listed in Table 6 and 7.
@@
Per-frame Tokenizer Input resolution $256\times 320$
Layers per block $2$
@@
Token number $N$ $16\times 20=320$
@@
Segment length $8$
Optimizer AdamW [35]
@@
Compressive Tokenizer [61] Channels $[128,256,256,512]$
@@
Context token number $N$ $32\times 40=1280$
@@
Number of sampled frames $7$
@@
We use the RT-1 dataset released from Open X-Embodiment [41], which follows the Apache license and contains 87,212 trajectories with $256\times 320$ visual observations and $13$-dimensional actions. Following Wu et al. [61], 99% of trajectories are used as the training set and 1% are left as the test set. During testing, a fixed segment from each trajectory is used for prediction.
@@
For single-step and multi-step prediction settings, we train a per-frame tokenizer and a compressive tokenizer [61] respectively on RT-1 from scratch. The per-frame tokenizer is essentially a VQGAN [10] with a convolutional encoder-decoder architecture, which tokenizes each frame $s_{t}$ in the trajectory independently into tokens $z_{t}\in[K]^{N}$ where $K$ is the codebook size. The compressive tokenizer is implemented as a conditional VQGAN with dual encoder-decoder pairs $\{(E_{c},D_{c}),(E_{p},D_{p})\}$. The context encoder $E_{c}$ first tokenizes a context frame $s_{c}$ independently into context tokens $z_{c}\in[K_{2}]^{N}$. Subsequently, each frame $s_{t}$ is tokenized by $E_{p}$, conditioned on the context encoder’s feature maps through cross attention, into tokens $z_{t}\in[K_{1}]^{n}$. Since the context features can capture rich shared information across the trajectory, the number of tokens per frame can be significantly reduced compared to independent tokenization ($n\ll N$). We refer to Wu et al. [61] for further details on the compressive tokenizer.
Architectural details of these tokenizers are provided in Table 6. Notably, we adopt finite scalar quantization (FSQ) [38] instead of vector quantization (VQ) [55] within our VQGANs, due to its superior codebook utilization.
@@
We adopt the standard architecture of LLaMA [54], instantiated as smaller models with 138M parameters, matching GPT-2 small [42]. Separate Transformers are pre-trained from scratch for single-step and multi-step prediction, respectively. During multi-step prediction training, context frames are sampled from earlier frames preceding the trajectory segment to prevent information leakage for prediction. Architecture and training are detailed in Table 7.
@@
Pre-training Training steps $9.9\times 10^{5}$ for single-step prediction
@@
Optimizer AdamW [35]
@@
Segment length $5$ for single-step prediction
$8$ for multi-step prediction
@@
RLVR Sequence length $1654$ for single-step prediction
@@
Optimizer AdamW [35]
@@
Learning rate $51\times 10^{-}$ 5
@@ A.5 Real2Sim Policy Evaluation
We aim to evaluate popular open-source generalist policies for the Google Robot, including a series of RT-1[5] checkpoints at different training stages: RT-1 trained to convergence (RT-1 (Converged)), RT-1 at 15% of training steps (RT-1 (15%)), and RT-1 at the beginning of training (RT-1 (Begin)), as well as RT-1-X [41]. SIMPLER [32] is a collection of simulated environments for manipulation policy evaluation on common real robot setups, including different methods to align with real-world environments: "SIMPLER-Visual Matching" (short as SIMPLER-VM in our paper) and "SIMPLER-Variant Aggregation" (short as SIMPLER-VA). We compare our world models against SIMPLER as the representation for policy evaluation methods based on physical simulation, which requires substantial human effort to develop task-specific environments.
@@
For each task, we randomly sample 30 initial frames from the real-world RT-1 dataset to serve as initial observations for world model simulations. Trajectories are then generated iteratively: (1) the policy selects the next action based on the latest frame; (2) the world model predicts the subsequent frame given the action. Following SIMPLER [32], we limit each trajectory to a maximum of 112 frames. Specifically, we use our trained world models for multi-step prediction, able to predict seven future frames conditioned on the current and initial context frame. To predict long trajectories, we apply a sliding window approach, where the last predicted frame becomes the input first frame in the next generation round. To conclude, Our evaluation covers two world models (the base model and RLVR-World), four policies (RT-1 (Begin), RT-1 (15%), RT-1 (Converged), and RT-1-X) [5, 41], six tasks (open/close top/middle/bottom drawer), and 30 trajectories per task, resulting in a total of $2\times 4\times 6\times 30=1440$ generated trajectories.
@@
Open Drawer Real 0.0% 51.9% 70.4% 81.5%
@@
Close Drawer Real 0.0% 74.1% 88.9% 92.6%
@@ Appendix B Computational Resources
The SFT phase uses 4$\times$ 80G A100 GPUs for 6.5 hours of training. The RLVR phase is conducted on 8$\times$ 80G A100 GPUs over 22.5 hours. Both SFT and RLVR are implemented using the verl framework [^8] https://github.com/volcengine/verl, following the Apache License. [50].
@@
SFT is performed on 8$\times$ 40G A100 GPUs over 17 hours, while the RLVR training is conducted on 8$\times$ 80G H100 GPUs for 25 hours. Both SFT and RLVR are implemented using the verl framework.
@@
All experiments in this domain are conducted on a 40G A100 GPU cluster. For single-step prediction, we pre-train the tokenizer using approximately 360 GPU hours. Transformer pre-training takes 530 GPU hours, and RLVR post-training for 200 steps requires 3.5 hours with 4$\times$ 40G A100 GPUs. For multi-step prediction, we pre-train the tokenizer with 480 GPU hours. Transformer pre-training takes 500 GPU hours, and RLVR post-training for 200 steps requires 10 hours with 4$\times$ 40G A100. Both settings implement tokenizer and transformer pre-training with accelerate [^9] https://github.com/huggingface/accelerate, following the Apache License. and RLVR post-training using verl.
@@ Appendix D Prompt Examples
In this section, we present detailed prompts used in our LLM experiments to offer a more intuitive understanding of language world models. These prompts are either directly taken from or adapted from prior work [57, 8].
@@
  • Web page state prediction and web agents : We present the prompts used in our study in orange boxes. Specifically, we include four types of prompts: Web State Prediction Prompt Example, Top-10 Salient Change Extraction Prompt, Next-State Change Summarization Prompt, and Web State Evaluation Prompt Example. The web state prediction prompt and next-change summarization prompt are modified versions of the original prompts proposed by WMA [8], while the top-10 salient change extraction prompt and value prediction (evaluation) prompt are used unchanged, adopted directly from WMA.
@@
[p] Text Game: State Difference Prediction Prompt Expample You are a simulator of a text game. Read the task description and the current environment observation description. Given the current game state in JSON, you need to decide the new game state after taking an action. Your response should be in the JSON format. It should have three keys: ‘modified’, ‘removed’, and ‘score’. The ‘modified’ key stores a list of all the object states that are added or changed after taking the action. Keep it an empty list if no object is added or modified. The ‘removed’ key stores a list of uuids of the objects that are removed. Keep it an empty list if no object is removed. The ‘score’ key stores a dictionary with three keys: ‘score’ is the current game score, ‘gameOver’ is a boolean of whether the game is over, and ‘gameWon’ is a boolean of whether the agent won the game. If a player earns a score or wins/loses the game, you should reflect that change in the dictionary saved under the ‘score’ key. Otherwise, you should set value of the ‘score’ key to an empty dictionary.Note that while game states can be changed by actions, some game states may change over the time, which is described in the tick function of each object class. Note that while game states can be changed by actions, some game states may change over the time, which is described in the tick function of each object class. Here are two examples of both cases. Both examples are from the same example game. Example game task description: Your task is to wash the dirty dishes. Here are the descriptions of all game objects properties in the example game: { OBJECT_RULES} Here are descriptions of all game actions in the example game: { ACTION_RULES} Here is a description of the score function of the example game: { SCORE_RULES} In the first example, the game state is changed by an action: Current observation: { GAME_OBSERVATION} Here is the game state: { GAME_STATE} The action to take is put dirty plate (ID: 5) in mug (ID: 6) The expected response is: { GAME_STATE_DIFFERENCE} In the second example from the same example game, the game state is changed over the time. Note that while in this example the game state is changed by time only, it is possible that a game state is changed by both an action and time. Current observation: {Example_2 observation} Here is the game state: { GAME_STATE} The action to take is eat dishwasher (ID: 2) with dirty plate (ID: 5) The expected response is: { GAME_STATE_DIFFERENCE} Here is the game that you need to simulate: Task Description: Your task is to boil water. Here are the descriptions of all game objects properties: { OBJECT_RULES} Here are the descriptions of all game actions: { ACTION_RULES} Here is a description of the score function of the game: { SCORE_RULES} Current observation: { GAME_OBSERVATION} Here is the game state: { GAME_STATE} The current game UUID base is 12 The action to take is: look
@@
[p] Text Game: Game State Example {‘game_state’: [{‘name’: ’agent (ID: 0)’, ‘uuid’: 0, ‘type’: ‘Agent’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’}, ‘contains’: [’plate (ID: 5)’, ’mug (ID: 6)’, ’knife (ID: 7)’]}, {‘name’: ’plate (ID: 5)’, ‘uuid’: 5, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘on’, ‘dishType’: ‘plate’, ‘isDirty’: True, ‘foodMessName’: ‘orange’}, ‘contains’: []}, {‘name’: ’mug (ID: 6)’, ‘uuid’: 6, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘mug’, ‘isDirty’: True, ‘foodMessName’: ‘sandwhich’}, ‘contains’: []}, {‘name’: ’knife (ID: 7)’, ‘uuid’: 7, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘knife’, ‘isDirty’: True, ‘foodMessName’: ’apple (ID: 11)’}, ‘contains’: []}, {‘name’: ’dishwasher (ID: 2)’, ‘uuid’: 2, ‘type’: ‘DishWasher’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: False, ‘isOpenable’: True, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘isDevice’: True, ‘isActivatable’: True, ‘isOn’: False, ‘cycleStage’: 0, ‘finishedCycle’: False}, ‘contains’: [’cup (ID: 4)’]}, {‘name’: ’cup (ID: 4)’, ‘uuid’: 4, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘cup’, ‘isDirty’: True, ‘foodMessName’: ’peanut butter’}, ‘contains’: []}, {‘name’: ’bottle of dish soap (ID: 3)’, ‘uuid’: 3, ‘type’: ‘DishSoapBottle’, ‘properties’: {‘isContainer’: False, ‘isMoveable’: True, ‘isDevice’: True, ‘isActivatable’: True, ‘isOn’: False}, ‘contains’: []}, {‘name’: ’glass (ID: 8)’, ‘uuid’: 8, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘glass’, ‘isDirty’: False}, ‘contains’: []}, {‘name’: ’bowl (ID: 9)’, ‘uuid’: 9, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘bowl’, ‘isDirty’: False}, ‘contains’: []}, {‘name’: ’banana (ID: 10)’, ‘uuid’: 10, ‘type’: ‘Food’, ‘properties’: {‘isContainer’: False, ‘isMoveable’: True, ‘isFood’: True}, ‘contains’: []}, {‘score’: -1, ‘gameOver’: False, ‘gameWon’: False}]}
@@
[p] Text Game: Game State Difference Example {‘modified’: [{‘name’: ’agent (ID: 0)’, ‘uuid’: 0, ‘type’: ‘Agent’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’}, ‘contains’: [’mug (ID: 6)’, ’knife (ID: 7)’]}, {‘name’: ’mug (ID: 6)’, ‘uuid’: 6, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘mug’, ‘isDirty’: True, ‘foodMessName’: ‘sandwhich’}, ‘contains’: [’plate (ID: 5)’]}], ‘removed’: [], ‘score’: {}}

### v2

@@ A.3 Model Predictive Control for Web Agents
Following the method introduced in WMA chae2024web , we use a policy model to sample 20 candidate actions and select the three most frequently sampled for further evaluation. For each selected action, our trained world model performs next-state prediction. A summarization model is then used to (1) identify the top 10 most salient state changes in the predicted next state, and (2) generate a natural language summary describing those changes. This transition summary, along with the task objective, is provided to a value model, which assigns a score between 1 and 5. This scoring query is repeated 20 times, and the final score for each action is computed as the average of these 20 responses. The action with the highest average score is selected for execution. Except for the world model, all models, including policy, summarization, and value, are implemented by prompting DeepSeek-V3 with top-$p$ sampling ($p=0.95$). All prompts used in model predictive control are detailed in Appendix D.
@@ A.4 Robot Manipulation Trajectory Prediction
A.4.1 RT-1
Hyperparameters of architectures and training process for robot manipulation trajectory prediction are listed in Table 7 and 8.
@@ A.4.1 RT-1
Per-frame
Tokenizer
Input resolution $256\times 320$
Layers per block 22
@@
Token number NN $16\times 20=320$
@@
Segment length 88
Optimizer AdamW loshchilov2017decoupled
@@
Compressive
Tokenizer wu2024ivideogpt
Channels $[128,256,256,512]$
@@
Context token number NN $32\times 40=1280$
@@
Number of sampled frames 77
@@
We use the RT-1 dataset released from Open X-Embodiment padalkar2023open , which follows the Apache license and contains 87,212 trajectories with $256\times 320$ visual observations and $13$-dimensional actions. Following Wu et al. wu2024ivideogpt , 99% of trajectories are used as the training set and 1% are left as the test set. During testing, a fixed segment from each trajectory is used for prediction.
@@
For single-step and multi-step prediction settings, we train a per-frame tokenizer and a compressive tokenizer wu2024ivideogpt respectively on RT-1 from scratch. The per-frame tokenizer is essentially a VQGAN esser2021taming with a convolutional encoder-decoder architecture, which tokenizes each frame $s_{t}$ in the trajectory independently into tokens $z_{t}\in[K]^{N}$ where $K$ is the codebook size. The compressive tokenizer is implemented as a conditional VQGAN with dual encoder-decoder pairs $\{(E_{c},D_{c}),(E_{p},D_{p})\}$. The context encoder $E_{c}$ first tokenizes a context frame $s_{c}$ independently into context tokens $z_{c}\in[K_{2}]^{N}$. Subsequently, each frame $s_{t}$ is tokenized by $E_{p}$, conditioned on the context encoder’s feature maps through cross attention, into tokens $z_{t}\in[K_{1}]^{n}$. Since the context features can capture rich shared information across the trajectory, the number of tokens per frame can be significantly reduced compared to independent tokenization ($n\ll N$). We refer to Wu et al. wu2024ivideogpt for further details on the compressive tokenizer.
Architectural details of these tokenizers are provided in Table 7. Notably, we adopt finite scalar quantization (FSQ) mentzer2023finite instead of vector quantization (VQ) van2017neural within our VQGANs, due to its superior codebook utilization.
@@
We adopt the standard architecture of LLaMA touvron2023llama , instantiated as smaller models with 138M parameters, matching GPT-2 small radford2019language . Separate Transformers are pre-trained from scratch for single-step and multi-step prediction, respectively. During multi-step prediction training, context frames are sampled from earlier frames preceding the trajectory segment to prevent information leakage for prediction. Architecture and training are detailed in Table 8.
@@
Pre-training
Training steps $9.9\times 10^{5}$ for single-step prediction
@@
Optimizer AdamW loshchilov2017decoupled
@@
Segment length 55 for single-step prediction
88 for multi-step prediction
@@
RLVR
Sequence length $1654$ for single-step prediction
@@
Optimizer AdamW loshchilov2017decoupled
@@
Learning rate $5\times 10^{-5}$
@@
A.4.2 PushT, Rope and Granular
Datasets.
We use the datasets provided by DINO-WM zhou2024dino . PushT is an environment introduced by Chi et al. chi2023diffusion that features a pusher agent interacting with a T-shaped block. The collected dataset contains 18,685 training trajectories and 21 test trajectories. Each step has a frame resolution of $224\times 224$ and includes proprioceptive information. The task of the dataset is to predict the next step based on the previous three steps. Rope and granular manipulation tasks are introduced by Zhang et al. zhang2024adaptigraph , simulated with Nvidia Flex. Each environment collects 900 training trajectories and 100 testing trajectories, in the resolution of $224\times 224$. The task of these two datasets is to predict the next step based on the previous one step.
Implementation details.
Unless otherwise specified, we adopt the same hyperparameters as in Appendix A.4.1 for RT-1. For PushT, proprioceptive information is quantized into uniform bins, consistent with the treatment of action values, and its mean squared error is included in the RLVR reward. For Rope and Granular, given the limited amount of training data, we employ a smaller model with 6 Transformer layers. A group size of $32$ is used in GRPO for all three datasets.
Learning curves.
The learning curves of our models on three datasets are shown in Table 9. Across all datasets, RLVR consistently improves upon our base model. Notably, when training with next-token prediction on smaller datasets (Rope and Granular), the models are prone to overfitting and require early stopping. In contrast, RLVR continues to significantly improve prediction performance on the validation set, demonstrating greater robustness and generalization.
Table 9: Learning curves of video world models on PushT, Rope, and Granular datasets. LPIPS scores are scaled by 100.
Pre-train RLVR
Training steps $1\times 10^{5}$ $2\times 10^{5}$ $3\times 10^{5}$ $60$ $100$ $200$ $300$ $400$
[…]
Pre-train RLVR
Training steps $2\times 10^{4}$ $4\times 10^{4}$ $6\times 10^{4}$ $7\times 10^{4}$ $100$ $200$ $400$ $720$
[…]
Pre-train RLVR
Training steps $1\times 10^{4}$ $2\times 10^{4}$ $3\times 10^{4}$ $40$ $100$ $200$ $300$ $500$
[…]
Extended analysis.
Our model underperforms DINO-WM on the Rope dataset. We hypothesize that this is due to our model being prone to overfitting on this relatively small dataset. In contrast, recurrent latent state models that leverage large-scale pretrained visual encoders demonstrate greater robustness. To validate this hypothesis, we explore training a base model jointly on the Rope and Granular datasets (both collected from PyFlex simulators), followed by fine-tuning with RLVR on each individual dataset. The results are shown in Table 11. We observe that the LPIPS on Rope improves from 0.0208 to 0.0165. Looking forward, we believe the potential of RLVR can be further unlocked when applied to large-scale pre-trained world models, mirroring the remarkable success of LLMs.
We further investigate whether improvements in one domain can benefit the other. As shown in the Table 11, we observe that when the model is RL-trained exclusively on one dataset, performance on the other can improve slightly. This result aligns with discoveries in the LLM field that improvements in one domain, such as math, can generalize to other domains. However, we also find that performance on the held-out dataset later degrades due to overfitting. We attribute this to the limitations of the base model, which is not pre-trained on large-scale, diverse data and thus lacks truly generalizable representations. We hope these preliminary results can encourage the community to further explore RLVR in the context of large-scale world models.
Table 10: Performance of video world models jointly trained on deformable object manipulation tasks. LPIPS and SSIM scores are scaled by 100.
Rope Granular
[…]
Base
(individually trained)
3.03 97.86 3.14 94.79
RLVR-World
(then individually tuned)
2.08 98.14 2.42 95.42
Base
(jointly trained)
2.59 97.95 3.41 94.54
RLVR-World
(then individually tuned)
1.65 98.29 2.44 95.40
Table 11: Cross-task evaluation of video world models jointly trained on deformable object manipulation tasks. LPIPS and SSIM scores are scaled by 100.
Rope Granular
[…]
Base
(jointly trained)
2.59 97.95 3.41 94.54
RLVR-World
(tuned on Rope)
1.65 98.29 3.35 94.60
RLVR-World
(tuned on Granular)
2.54 97.97 2.44 95.40
Qualitative analysis.
Showcases of our RLVR-World models are presented in Figure 10.
Figure 10: Qualitative analysis: video prediction on PushT, Rope, and Granular.
Model-predictive control.
We strictly follow the same planner configuration as DINO-WM, including goal construction, number of samples, and optimization steps. Since our model predicts discrete tokens and raw pixels without a compact latent space, we use the publicly available DINOv2 encoder to embed predicted frames and compare them to the goal observation during planning. We find that the DINO latent space is more effective for planning than pixel-level MSE or LPIPS, consistent with the findings in DINO-WM. We emphasize that our MPC experiments are intended to compare world models, and the choice of distance metric is orthogonal to our main contributions.
We do not conduct experiments on Rope and Granular due to the lack of MPC configurations and details in the official DINO-WM code, making it difficult to replicate the setup.
The results in Table 12 show that RLVR also enhances the control performance of our base model, achieving results comparable to DINO-WM.
Table 12: Model-predictive control with different world models on PushT.
Model IRIS DreamerV3 TD-MPC2
CEM w/
DINO-WM
CEM w/
Our Base
CEM w/
Our RLVR-World
PushT
Success Rate
0.32 0.30 0.00 0.86 0.80 0.86
@@ A.5 Real2Sim Policy Evaluation
We aim to evaluate popular open-source generalist policies for the Google Robot, including a series of RT-1brohan2022rt checkpoints at different training stages: RT-1 trained to convergence (RT-1 (Converged)), RT-1 at 15% of training steps (RT-1 (15%)), and RT-1 at the beginning of training (RT-1 (Begin)), as well as RT-1-X padalkar2023open . SIMPLER li2024evaluating is a collection of simulated environments for manipulation policy evaluation on common real robot setups, including different methods to align with real-world environments: "SIMPLER-Visual Matching" (short as SIMPLER-VM in our paper) and "SIMPLER-Variant Aggregation" (short as SIMPLER-VA). We compare our world models against SIMPLER as the representation for policy evaluation methods based on physical simulation, which requires substantial human effort to develop task-specific environments.
@@
For each task, we randomly sample 30 initial frames from the real-world RT-1 dataset to serve as initial observations for world model simulations. Trajectories are then generated iteratively: (1) the policy selects the next action based on the latest frame; (2) the world model predicts the subsequent frame given the action. Following SIMPLER li2024evaluating , we limit each trajectory to a maximum of 112 frames. Specifically, we use our trained world models for multi-step prediction, able to predict seven future frames conditioned on the current and initial context frame. To predict long trajectories, we apply a sliding window approach, where the last predicted frame becomes the input first frame in the next generation round. To conclude, Our evaluation covers two world models (the base model and RLVR-World), four policies (RT-1 (Begin), RT-1 (15%), RT-1 (Converged), and RT-1-X) brohan2022rt ; padalkar2023open , six tasks (open/close top/middle/bottom drawer), and 30 trajectories per task, resulting in a total of $2\times 4\times 6\times 30=1440$ generated trajectories.
@@
Open
Drawer
Real 0.0% 51.9% 70.4% 81.5%
@@
Close
Drawer
Real 0.0% 74.1% 88.9% 92.6%
@@ Appendix B Computational Resources
[…]
@@
[…]
@@
All experiments in this domain are conducted on a 40G A100 GPU cluster. For single-step prediction, we pre-train the tokenizer using approximately 360 GPU hours. Transformer pre-training takes 530 GPU hours, and RLVR post-training for 200 steps requires 3.5 hours with $\times 4$0G A100 GPUs. For multi-step prediction, we pre-train the tokenizer with 480 GPU hours. Transformer pre-training takes 500 GPU hours, and RLVR post-training for 200 steps requires 10 hours with $\times 4$0G A100. Both settings implement tokenizer and transformer pre-training with accelerate [^9] https://github.com/huggingface/accelerate, following the Apache License. and RLVR post-training using verl.
@@ Appendix D Prompt Examples
In this section, we present detailed prompts used in our LLM experiments to offer a more intuitive understanding of language world models. These prompts are either directly taken from or adapted from prior work wang2024can ; chae2024web .
@@
  • Web page state prediction and web agents : We present the prompts used in our study in orange boxes. Specifically, we include four types of prompts: Web State Prediction Prompt Example, Top-10 Salient Change Extraction Prompt, Next-State Change Summarization Prompt, and Web State Evaluation Prompt Example. The web state prediction prompt and next-change summarization prompt are modified versions of the original prompts proposed by WMA chae2024web , while the top-10 salient change extraction prompt and value prediction (evaluation) prompt are used unchanged, adopted directly from WMA.
@@
Text Game: State Difference Prediction Prompt Expample You are a simulator of a text game. Read the task description and the current environment observation description. Given the current game state in JSON, you need to decide the new game state after taking an action. Your response should be in the JSON format. It should have three keys: ‘modified’, ‘removed’, and ‘score’. The ‘modified’ key stores a list of all the object states that are added or changed after taking the action. Keep it an empty list if no object is added or modified. The ‘removed’ key stores a list of uuids of the objects that are removed. Keep it an empty list if no object is removed. The ‘score’ key stores a dictionary with three keys: ‘score’ is the current game score, ‘gameOver’ is a boolean of whether the game is over, and ‘gameWon’ is a boolean of whether the agent won the game. If a player earns a score or wins/loses the game, you should reflect that change in the dictionary saved under the ‘score’ key. Otherwise, you should set value of the ‘score’ key to an empty dictionary.Note that while game states can be changed by actions, some game states may change over the time, which is described in the tick function of each object class. Note that while game states can be changed by actions, some game states may change over the time, which is described in the tick function of each object class. Here are two examples of both cases. Both examples are from the same example game. Example game task description: Your task is to wash the dirty dishes. Here are the descriptions of all game objects properties in the example game: { OBJECT_RULES} Here are descriptions of all game actions in the example game: { ACTION_RULES} Here is a description of the score function of the example game: { SCORE_RULES} In the first example, the game state is changed by an action: Current observation: { GAME_OBSERVATION} Here is the game state: { GAME_STATE} The action to take is put dirty plate (ID: 5) in mug (ID: 6) The expected response is: { GAME_STATE_DIFFERENCE} In the second example from the same example game, the game state is changed over the time. Note that while in this example the game state is changed by time only, it is possible that a game state is changed by both an action and time. Current observation: {Example_2 observation} Here is the game state: { GAME_STATE} The action to take is eat dishwasher (ID: 2) with dirty plate (ID: 5) The expected response is: { GAME_STATE_DIFFERENCE} Here is the game that you need to simulate: Task Description: Your task is to boil water. Here are the descriptions of all game objects properties: { OBJECT_RULES} Here are the descriptions of all game actions: { ACTION_RULES} Here is a description of the score function of the game: { SCORE_RULES} Current observation: { GAME_OBSERVATION} Here is the game state: { GAME_STATE} The current game UUID base is 12 The action to take is: look
@@
Text Game: Game State Example {‘game_state’: [{‘name’: ’agent (ID: 0)’, ‘uuid’: 0, ‘type’: ‘Agent’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’}, ‘contains’: [’plate (ID: 5)’, ’mug (ID: 6)’, ’knife (ID: 7)’]}, {‘name’: ’plate (ID: 5)’, ‘uuid’: 5, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘on’, ‘dishType’: ‘plate’, ‘isDirty’: True, ‘foodMessName’: ‘orange’}, ‘contains’: []}, {‘name’: ’mug (ID: 6)’, ‘uuid’: 6, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘mug’, ‘isDirty’: True, ‘foodMessName’: ‘sandwhich’}, ‘contains’: []}, {‘name’: ’knife (ID: 7)’, ‘uuid’: 7, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘knife’, ‘isDirty’: True, ‘foodMessName’: ’apple (ID: 11)’}, ‘contains’: []}, {‘name’: ’dishwasher (ID: 2)’, ‘uuid’: 2, ‘type’: ‘DishWasher’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: False, ‘isOpenable’: True, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘isDevice’: True, ‘isActivatable’: True, ‘isOn’: False, ‘cycleStage’: 0, ‘finishedCycle’: False}, ‘contains’: [’cup (ID: 4)’]}, {‘name’: ’cup (ID: 4)’, ‘uuid’: 4, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘cup’, ‘isDirty’: True, ‘foodMessName’: ’peanut butter’}, ‘contains’: []}, {‘name’: ’bottle of dish soap (ID: 3)’, ‘uuid’: 3, ‘type’: ‘DishSoapBottle’, ‘properties’: {‘isContainer’: False, ‘isMoveable’: True, ‘isDevice’: True, ‘isActivatable’: True, ‘isOn’: False}, ‘contains’: []}, {‘name’: ’glass (ID: 8)’, ‘uuid’: 8, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘glass’, ‘isDirty’: False}, ‘contains’: []}, {‘name’: ’bowl (ID: 9)’, ‘uuid’: 9, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘bowl’, ‘isDirty’: False}, ‘contains’: []}, {‘name’: ’banana (ID: 10)’, ‘uuid’: 10, ‘type’: ‘Food’, ‘properties’: {‘isContainer’: False, ‘isMoveable’: True, ‘isFood’: True}, ‘contains’: []}, {‘score’: -1, ‘gameOver’: False, ‘gameWon’: False}]}
@@
Text Game: Game State Difference Example {‘modified’: [{‘name’: ’agent (ID: 0)’, ‘uuid’: 0, ‘type’: ‘Agent’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’}, ‘contains’: [’mug (ID: 6)’, ’knife (ID: 7)’]}, {‘name’: ’mug (ID: 6)’, ‘uuid’: 6, ‘type’: ‘Dish’, ‘properties’: {‘isContainer’: True, ‘isMoveable’: True, ‘isOpenable’: False, ‘isOpen’: True, ‘containerPrefix’: ‘in’, ‘dishType’: ‘mug’, ‘isDirty’: True, ‘foodMessName’: ‘sandwhich’}, ‘contains’: [’plate (ID: 5)’]}], ‘removed’: [], ‘score’: {}}